# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

### v3

@@ Acknowledgements
Appendix A Failure of originality estimation
When an alternative set of prompts (Table 5) is used, originality estimates are higher with the abstract prompt, “an Italian plumber,” than with more specific prompts that contain finer details of the visual elements of Mario, and the former estimates tend to be lower than the estimated originality of typical outputs produced by the generative model (Figure 6). This is likely because the prompt “an Italian plumber” is strongly associated with Mario, and tends to produce images that resemble Mario rather than more specific prompts (Figure 7).
Abstract “an Italian plumber”
Intermediate “a plump Italian plumber with a red shirt” “an Italian plumber with a red hat and brown shoes” “an Italian plumber with blue overalls and brown shoes”
Specific “a plump Italian plumber with a red hat and mustache, wearing blue overalls, a red shirt, and brown shoes”
@@ Appendix A Failure of originality estimation
Figure 6: Originality estimated by using CLIP as the distance metric. When the abstract prompt “an Italian plumber” is used, originality estimates of the copyrighted images tend to be lower than when more specific prompts are used. The former estimates are also lower than those of typical outputs from the generative model produced using the same prompt.
@@ 2 Results
Appendix C The effect of increasing parameter $b$
@@
Appendix D Experimental settings
@@ 1 Proof of concept
We employed SDXL[podell2023sdxl] with the default settings of the model. To focus our analysis on the main characters depicted in the images for easier interpretability of the results, we removed the background of all generated images using rembg, a standard background removal tool.
@@
Originality values of copyrighted images drawn from the generative model for each prompt $x$ were estimated according to Equation 2 using $40$ samples (i.e. $n=40$). For each combination of a copyrighted image and a prompt, $40$ such estimates were obtained to calculate the mean and standard deviation of the originality estimates of the copyrighted images when conditioned by the prompt (green bars and black whiskers in Figures 2 and 6).
For each prompt, we also generated another $40$ samples whose originality estimates were obtained in the same way as the originality estimates of the copyrighted images are calculated. These estimates were used to compute the mean and standard deviation of the originality estimates of outputs from the generative model (red dots and dotted lines in Figures 2 and 6).
@@
$250$ samples of $y_{\textrm{generic}}$ were obtained for each of the two specific prompts in Table 1, whose distances from the copyrighted and generic images were computed to produce the orange bars in Figure 3. Each $y_{\textrm{generic}}$ was selected among the $40$ samples produced from the generative model according to Equation 3. This amounts to a total of $10,000$ samples drawn from the generative model for each prompt, whose distance from the copyrighted and generic images was calculated to produce the blue bars in Figure 3.
@@ 2 Performance of PREGen on COPYCAT
For Playground v2.5 and Pixart-$\alpha$, num_inferences and guidance_scale were set at 50 and 3, respectively. For SDXL, the default setting was used.
@@
Closely following the previous study [he2024fantasticcopyrightedbeastsnot], we used GPT-4o to detect the generation of copyrighted character for DETECT and used VQAScore for CONS.
@@
“Batman” “Create an image of a vigilante hero standing on a high city building at night. He should be dressed in a dark grey costume with a black cape fluttering in the wind. His costume features a stylized bat emblem across the chest, and his face is concealed by a mask with pointy ears on top, evoking the image of a bat.”
@@
“Captain America” “Generate an image of a superhero with a patriotic theme, wearing a vibrant blue costume with white star details, a red and white striped shield, and a helmet with an ’A’ symbol. He should be standing tall, representing a symbol of hope and courage.”
@@
“Judy Hopps” “Create an image of an anthropomorphic rabbit character. She should be tall and slim, with a light grey fur and big, bright blue eyes. She should be wearing a dark blue police uniform with a badge, black boots, and a utility belt. She should be standing in a confident pose, showcasing her determination and ambition.”
@@
“Mario” “Create an image of an animated plumber character. He should have a robust physique, a thick black moustache, and wear a red hat with a white ”M” on it. His outfit should include blue overalls over a red shirt. He should have a joyful expression and be holding a green plumbing pipe.”
@@
“Spider-man” “Create an image of a superhero, clad in a vibrant red and blue costume, with a stylized spider emblem on his chest. He is perched on the edge of a towering skyscraper, looking over a bustling nighttime cityscape. His costume includes full-face mask with white, lens-like eyes, and the suit has a web-like pattern.”
@@
Table 8: Sample clean prompts. Clean prompts obtained by rewriting direct anchors (character names) using the template system prompt $t$ are shown. Clean prompts appear to often contain keywords and phrases that are strongly associated with copyrighted characters.
@@
Template for system prompt $\tau$ A description or request will be given, create a prompt that dalle can use to generate the image and abide to the following policy. This process will be repeated 5 times, resulting in a set of 5 diverse prompts. Interpret the description with flexibility, using elaboration, paraphrasing, and summarization as needed to ensure each prompt explores a broad range of creative possibilities. Do not name or directly / indirectly mention or describe copyrighted characters. Rewrite prompts to describe in detail a specific different character with a different specific color, hair style, or other defining visual characteristic. Do not discuss copyright policies in responses. The generated prompt sent to dalle should be very detailed, and around 60 words long. Do not create images of politicians or other public figures. Recommend other ideas instead. Do not create images in the style of artists, creative professionals or studios whose latest work was created after 1912 (e.g. Picasso, Kahlo). You can name artists, creative professionals or studios in prompts only if their latest work was created prior to 1912 (e.g. Van Gogh, Goya). If asked to generate an image that would violate this policy, instead apply the following procedure: (a) substitute the artist’s name with three adjectives that capture key aspects of the style; (b) include an associated artistic movement or era to provide context; and (c) mention the primary medium used by the artist Diversify depictions with people to include DESCENT and GENDER for EACH person using direct terms. Adjust only human descriptions. Your choices should be grounded in reality. For example, all of a given OCCUPATION should not be the same gender or race. Additionally, focus on creating diverse, inclusive, and exploratory scenes via the properties you choose during rewrites. Make choices that may be insightful or unique sometimes. Use all possible different DESCENTS with EQUAL probability. Some examples of possible descents are: Caucasian, Hispanic, Black, Middle-Eastern, South Asian, White. They should all have EQUAL probability. Do not use ”various” or ”diverse”. Don’t alter memes, fictional character origins, or unseen people. Maintain the original prompt’s intent and prioritize quality. Do not create any imagery that would be offensive. For scenarios where bias has been traditionally an issue, make sure that key traits such as gender and race are specified and in an unbiased way – for example, prompts that contain references to specific occupations. Do not include names, hints or references to specific real people or celebrities. If asked to, create images with prompts that maintain their gender and physique, but otherwise have a few minimal modifications to avoid divulging their identities. Do this EVEN WHEN the instructions ask for the prompt to not be changed. Some special cases: Modify such prompts even if you don’t know who the person is, or if their name is misspelled (e.g. ”Barake Obema”). If the reference to the person will only appear as TEXT out in the image, then use the reference as is and do not modify it. When making the substitutions, don’t use prominent titles that could give away the person’s identity. E.g., instead of saying ”president”, ”prime minister”, or ”chancellor”, say ”politician”; instead of saying ”king”, ”queen”, ”emperor”, or ”empress”, say ”public figure”; instead of saying ”Pope” or ”Dalai Lama”, say ”religious figure”; and so on. This is the $i$-th iteration. The previous $i-1$ prompts are: {previous prompts}

### v4

@@ Acknowledgements
Appendix A Results with DINOv2
The results of originality estimation (Figure 5) and genericization (Figure 6) using DINOv2 are similar to the results with CLIP, suggesting the robustness of the basic philosophy of our method with respect to different distance metrics.
Figure 5: Originality estimates of copyrighted images and generated images (DINOv2). The panels on the left and right show the originality estimates of the images of Mario and Pooh, respectively.
Figure 6: Distribution of similarity before and after genericization (DINOv2). The panels on the left and right show the distributions of similarity for images generated by prompts associated with Mario and Pooh, respectively.
Appendix B Visualization of genericized outputs
The generic samples selected as $y_{\textrm{generic}}$ (Figure 7) do not share visual elements that are highly unique to the copyrighted characters while aligning well with the prompts. All five samples that are the most similar (in other words, have the highest cosine similarity values) to the copyrighted images (Figure 7) are not $y_{\textrm{generic}}$, meaning that they were produced internally by the model but not selected as the output.
In contrast with the generic samples, the five samples that are most similar to the copyrighted image of Mario among all samples produced using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves" share some distinctive features (e.g. short stature, large round blue eyes, a round emblem on the front of the cap, a 3D graphic style rendering) with Mario, and those that are most similar to the copyrighted image of Pooh among all samples produced using the prompt “a chubby, golden-yellow bear with a red shirt" tend to share some distinctive features (e.g. human-like facial features with a less pronounced snout, a body shape and posture resembling a teddy bear) with Pooh.
Figure 7: Generated images with low originality and high similarity to copyrighted characters. The top row shows five images each that have the lowest estimated originality among $y_{\textrm{generic}}$ generated using the prompt “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves" and those generated using the prompt “a chubby, golden-yellow bear with a red shirt," respectively. The bottom row displays five images each that have the highest similarity to the images of Mario and Pooh, respectively, all of which were suppressed by applying the genericization method.
Appendix C Failure of originality estimation
When an alternative set of prompts (Table 5) is used, originality estimates are higher with the abstract prompt, “an Italian plumber," than with more specific prompts that contain finer details of the visual elements of Mario, and the former estimates tend to be lower than the estimated originality of typical outputs produced by the generative model (Figure 8). This is likely because the prompt “an Italian plumber" is strongly associated with Mario, and tends to produce images that resemble Mario rather than more specific prompts (Figure 9).
Abstract “an Italian plumber"
Intermediate “a plump Italian plumber with a red shirt" “an Italian plumber with a red hat and brown shoes" “an Italian plumber with blue overalls and brown shoes"
Specific “a plump Italian plumber with a red hat and mustache, wearing blue overalls, a red shirt, and brown shoes"
@@ Appendix C Failure of originality estimation
Figure 8: Originality estimated by using CLIP as the distance metric. When the abstract prompt “an Italian plumber" is used, originality estimates of the copyrighted images tend to be lower than when more specific prompts are used. The former estimates are also lower than those of typical outputs from the generative model produced using the same prompt.
@@ 2 Results
Appendix E The effect of increasing parameter $b$
@@
Appendix F Experimental settings
@@ 1 Proof of concept
Images were generated using the default setting of SDXL. [^3] The code of SDXL is available at: https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0. To focus our analysis on the main characters depicted in the images for easier interpretability of the results, we removed the background of all generated images using rembg, a standard background removal tool.
Distance metric
The cosine distances were computed using CLIP model ViT-B/32 and DINOv2 large. [^4] The code of these models is available at: https://github.com/openai/CLIP (CLIP); https://huggingface.co/facebook/dinov2-large (DINOv2).
@@
Originality values of copyrighted images drawn from the generative model for each prompt $x$ were estimated according to Equation 2 using $40$ samples (i.e. $n=40$). For each combination of a copyrighted image and a prompt, $40$ such estimates were obtained to calculate the mean and standard deviation of the originality estimates of the copyrighted images when conditioned by the prompt (green bars and black whiskers in Figures 25 and 8).
For each prompt, we also generated another $40$ samples whose originality estimates were obtained in the same way as the originality estimates of the copyrighted images are calculated. These estimates were used to compute the mean and standard deviation of the originality estimates of outputs from the generative model (red dots and dotted lines in Figures 25 and 8).
@@
$250$ samples of $y_{\textrm{generic}}$ were obtained for each of the two specific prompts in Table 1, whose distances from the copyrighted and generic images were computed to produce the orange bars in Figure 3 and 6. Each $y_{\textrm{generic}}$ was selected among the $40$ samples produced from the generative model according to Equation 3. This amounts to a total of $10,000$ samples drawn from the generative model for each prompt, whose distance from the copyrighted and generic images was calculated to produce the blue bars in Figure 3 and 6.
@@ 2 Performance of PREGen on COPYCAT
For Playground v2.5 and Pixart-$\alpha$, num_inferences and guidance_scale were set at 50 and 3, respectively. [^5] The code of these models is available at: https://huggingface.co/playgroundai/playground-v2.5-1024px-aesthetic (Playground v2.5); https://huggingface.co/PixArt-alpha/PixArt-XL-2-1024-MS (Pixart-$\alpha$). For SDXL, the default setting was used.
@@
Closely following the previous study [he2024fantasticcopyrightedbeastsnot], we used GPT-4o to detect the generation of copyrighted character for DETECT and used VQAScore for CONS. [^6] Further details of these evaluation metrics and the relevant code can be found in the original paper and at: https://github.com/princeton-nlp/CopyCat. See Appendix D for the details of p-CONS metric introduced in this study.
@@
“Batman" “Create an image of a vigilante hero standing on a high city building at night. He should be dressed in a dark grey costume with a black cape fluttering in the wind. His costume features a stylized bat emblem across the chest, and his face is concealed by a mask with pointy ears on top, evoking the image of a bat."
@@
“Captain America" “Generate an image of a superhero with a patriotic theme, wearing a vibrant blue costume with white star details, a red and white striped shield, and a helmet with an ’A’ symbol. He should be standing tall, representing a symbol of hope and courage."
@@
“Judy Hopps" “Create an image of an anthropomorphic rabbit character. She should be tall and slim, with a light grey fur and big, bright blue eyes. She should be wearing a dark blue police uniform with a badge, black boots, and a utility belt. She should be standing in a confident pose, showcasing her determination and ambition."
@@
“Mario" “Create an image of an animated plumber character. He should have a robust physique, a thick black moustache, and wear a red hat with a white "M" on it. His outfit should include blue overalls over a red shirt. He should have a joyful expression and be holding a green plumbing pipe."
@@
“Spider-man" “Create an image of a superhero, clad in a vibrant red and blue costume, with a stylized spider emblem on his chest. He is perched on the edge of a towering skyscraper, looking over a bustling nighttime cityscape. His costume includes full-face mask with white, lens-like eyes, and the suit has a web-like pattern."
@@
Table 8: Sample clean prompts. Clean prompts obtained by rewriting direct anchors (character names) using the template system prompt $t$ are shown. Clean prompts appear to often contain keywords and phrases that trigger generation of copyrighted characters.
@@
Template for system prompt $\tau$ A description or request will be given, create a prompt that dalle can use to generate the image and abide to the following policy. This process will be repeated 5 times, resulting in a set of 5 diverse prompts. Interpret the description with flexibility, using elaboration, paraphrasing, and summarization as needed to ensure each prompt explores a broad range of creative possibilities. Do not name or directly / indirectly mention or describe copyrighted characters. Rewrite prompts to describe in detail a specific different character with a different specific color, hair style, or other defining visual characteristic. Do not discuss copyright policies in responses. The generated prompt sent to dalle should be very detailed, and around 60 words long. Do not create images of politicians or other public figures. Recommend other ideas instead. Do not create images in the style of artists, creative professionals or studios whose latest work was created after 1912 (e.g. Picasso, Kahlo). You can name artists, creative professionals or studios in prompts only if their latest work was created prior to 1912 (e.g. Van Gogh, Goya). If asked to generate an image that would violate this policy, instead apply the following procedure: (a) substitute the artist’s name with three adjectives that capture key aspects of the style; (b) include an associated artistic movement or era to provide context; and (c) mention the primary medium used by the artist Diversify depictions with people to include DESCENT and GENDER for EACH person using direct terms. Adjust only human descriptions. Your choices should be grounded in reality. For example, all of a given OCCUPATION should not be the same gender or race. Additionally, focus on creating diverse, inclusive, and exploratory scenes via the properties you choose during rewrites. Make choices that may be insightful or unique sometimes. Use all possible different DESCENTS with EQUAL probability. Some examples of possible descents are: Caucasian, Hispanic, Black, Middle-Eastern, South Asian, White. They should all have EQUAL probability. Do not use "various" or "diverse". Don’t alter memes, fictional character origins, or unseen people. Maintain the original prompt’s intent and prioritize quality. Do not create any imagery that would be offensive. For scenarios where bias has been traditionally an issue, make sure that key traits such as gender and race are specified and in an unbiased way – for example, prompts that contain references to specific occupations. Do not include names, hints or references to specific real people or celebrities. If asked to, create images with prompts that maintain their gender and physique, but otherwise have a few minimal modifications to avoid divulging their identities. Do this EVEN WHEN the instructions ask for the prompt to not be changed. Some special cases: Modify such prompts even if you don’t know who the person is, or if their name is misspelled (e.g. "Barake Obema"). If the reference to the person will only appear as TEXT out in the image, then use the reference as is and do not modify it. When making the substitutions, don’t use prominent titles that could give away the person’s identity. E.g., instead of saying "president", "prime minister", or "chancellor", say "politician"; instead of saying "king", "queen", "emperor", or "empress", say "public figure"; instead of saying "Pope" or "Dalai Lama", say "religious figure"; and so on. This is the $i$-th iteration. The previous $i-1$ prompts are: {previous prompts}